Use the following information to calculate the Cash Conversion Cycle for this company:

Year 1 Year 2
Annual Credit Sales 8
COGS 43
Accounts Receivable $2 4
Inventory $8 10
Inventory Purchases 47
Accounts payable $7 9

Round your answer to 2 decimals, for example 100.12.
Answer #1

Cash conversion cycle = Days inventory outstanding + Days sales outstanding + Days payable outstanding


Days inventory outstanding = (Average inventory/Cost of goods sold) * 365

Average inventory = (8+10+47)/2 = 32.5

Days inventory outstanding = (32.5/43) * 365 = 275.8721


Days sales outstanding = (Average accounts receivables / Total credit sales) * 365

Average accounts receivables = (2+4)/2 = 3

Days sales outstanding = (3/8) * 365 = 136.8750


Days payable outstanding = (Average accounts payable / cost of goods sold) * 365

Average accounts payable = (7+9)/2 = 8

Days payable outstanding = (8/43)*365 = 67.9070


Cash conversion cycle = 275.8721 + 136.8750 + 67.9070 = 480.65 days

FEEL FREE TO ASK DOUBTS.